OpenStudy (anonymous):

Two lines drawn at right angles, one being a tangent to \(y^2=4ax\) and the other to \(x^2=4by\). Find the locus of their point of intersection.

OpenStudy (sasogeek):

is this like one of those challenge questions?

OpenStudy (anonymous):

It ain't challenge question, I have already solved it but I need a better way.

OpenStudy (dumbcow):

sorry for not posting any work. here is what i came up with, Let c be any given x on the curve x^2 =4by Then the point of intersection of tangent line at c with perpendicular tangent line is (x*,y*): \[x^{*} = \frac{c^{2}(c-2a)}{2(c^{2}+4b^{2})}\] \[y^{*} = \frac{-c(a+2b^{2}c)}{2b(c^{2}+4b^{2})}\]